Recall that the variance of a random variable is defined as

Var[X]=E[(X−μ)2], where μ = E[X]. Use the properties of expectation to show that we can rewrite the variance of a random variable X as Var [X]=E[X^2]−(E[X])^2

Problem 3. (1 point) Recall that the variance of a random variable is defined as Var X-E(X-μ)21, where μ= E[X]. Use the properties of expectation to show that we can rewrite the variance of a random variable X as u hare i- ElX)L a
Answer #1

Var(X) =E(X-E(X))2 =E(X2+(E(X))2-2XE(X)) =E(X2)+E(((E(X))2-2*E(x*E(X))

=E(X2)+(E(X))2-2*E(X)*E(X) =E(X2)+(E(X))2-2*(E(X))2

Var(X)=E(X2)-(E(X))2